# Exploring Multilingual Concepts of Human Values in Large Language Models: Is Value Alignment Consistent, Transferable and Controllable across Languages?

### v2

@@ Abstract
Prior research has revealed that abstract concepts are linearly represented as directions in the representation space of LLMs, predominantly centered around English. In this paper, we extend this investigation to a multilingual context, with a specific focus on human values-related concepts (i.e., value concepts) due to their significance for AI safety. Through our comprehensive exploration covering 7 types of human values, 16 languages and 3 LLM series with distinct multilinguality, we first empirically confirm the presence of value concepts within LLMs in a multilingual format. Further analysis on the cross-lingual characteristics of these concepts reveals 3 traits arising from language resource disparities: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, all in terms of value concepts. Moreover, we validate the feasibility of cross-lingual control over value alignment capabilities of LLMs, leveraging the dominant language as a source language. Ultimately, recognizing the significant impact of LLMs’ multilinguality on our results, we consolidate our findings and provide prudent suggestions on the composition of multilingual data for LLMs pre-training.
Warning: This paper contains examples that can be upsetting or offensive.
@@
Shaoyang Xu1, Weilong Dong2, Zishan Guo2, Xinwei Wu2and Deyi Xiong21thanks:  Corresponding author. 1School of New Media and Communication, Tianjin University, Tianjin, China 2College of Intelligence and Computing, Tianjin University, Tianjin, China {syxu, willowd, guozishan, wuxw2021, dyxiong}@tju.edu.cn
@@ 1 Introduction
Recent years have witnessed the emergence of large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and LLaMA2 (Touvron et al., 2023). These LLMs have shown powerful capabilities in natural language understanding and generation (Guo et al., 2023; Bang et al., 2023; Jiao et al., 2023). However, alongside with their prowess, LLMs present potential risks. Research has demonstrated that LLMs can generate responses containing toxic, untruthful, biased, and even illegal content (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Thus, aligning LLMs with human values (i.e., value alignment) is necessary for unleashing their potential safely.
@@
However, existing studies on concept representations in LLMs have primarily focused on English, leaving multilingual concepts unexplored. Our work is the first to explore multilingual concepts in LLMs, emphasizing human values-related concepts to advance multilingual AI safety and utility. The primary research questions we aim to answer are as follows: (Q1) Do LLMs encode concepts representing human values in multiple languages? (Q2) To what extent are these concepts consistent and transferable across different languages? (Q3) Whether LLMs trained with different distributions of multilingual data exhibit distinct multilinguality in these concepts? (Q4) Is Value Alignment of LLMs Controllable across Languages? To address these questions, we propose a framework consisting of 5 essential components: extracting multilingual concept vectors from LLMs (§3.1) and evaluating their correlation with the corresponding concepts (concept recognition task in §3.2) to answer Q1; computing cross-lingual similarity of concept vectors (§3.3) and performing cross-lingual concept recognition (§3.4) to answer Q2 and Q3; and manipulating model behavior cross-lingually via concept vectors (§5) to answer Q4.
@@
Our analysis covers 7 concepts related to human values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness, given their significance for AI safety (Hendrycks et al., 2021; Bai et al., 2022; Askell et al., 2021; Touvron et al., 2023). To ensure the breadth and reliability of our findings, we have selected these 7 concepts for their diverse definitions and ethical attributes (Vida et al., 2023). Throughout this paper, we collectively refer to them as “value concepts” to reflect their diversity and keep consistent with existing AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). For comprehensive definitions, ethical backgrounds and examples of these value concepts, please refer to Appendix A.
@@
In addition to diverse human values, our experiments involve 16 languages [^1] We recognize that linguistic diversity can foster cultural variations, potentially resulting in diverse interpretations of the same value from different cultural backgrounds (Hershcovich et al., 2022; Hämmerl et al., 2023). For example, regarding deontology, some cultures prioritize individual responsibility while others emphasize social obligations (Cao et al., 2023; Hofstede, 1984). However, our work focuses on the multilingual representations of value concepts within LLMs and their universal cross-lingual patterns, leaving the exploration on cultural divergences in human values for our future research. and 3 LLM families with different multilinguality. Specifically, we categorize the multilinguality of these 3 LLM families based on language distributions in their pre-training data into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with respectively balanced multilinguality (i.e., BLOOMZ series). Appendix D provides detailed language distributions of their pre-training data.
@@ 2 Related Work
Representation Engineering (RepE) introduced by Zou et al. (2023a) extracts abstract concepts as vectors from LLMs using positive and negative samples that describe specific concepts. The effectiveness of these vectors has been validated across dimensions such as correlation and manipulation. Specifically, correlation experiments have assessed the predictive power of the extracted vectors to classify out-of-distribution data as positive or negative, while manipulation experiments have evaluated the vectors’ ability to control LLMs’ behavior by adding or subtracting them from the hidden states (Liu et al., 2023; Leong et al., 2023; Wang and Shu, 2023). While previous research has primarily focused on English, we pioneers the extension of RepE into a multilingual context. exploring multilingual concepts within LLMs through concept extraction, correlation, and manipulation experiments, all conducted in a multilingual or cross-lingual manner.
@@
Despite their remarkable capabilities, LLMs present potential risks (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023). Research has indicated that LLMs exhibit even greater vulnerability in multilingual scenarios, as evidenced by the multilingual jailbreak challenge (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). In this work, we explore the cross-lingual characteristics of safety-related concepts in LLMs’ representation spaces, shedding light on the mechanisms underlying multilingual vulnerability.
@@ 3.2 Recognizing Multilingual Concepts
Note that each layer has a recognition accuracy, using the concept vector of that layer. Generally, intermediate layers encode the most precise concepts. Unless specified otherwise, we select the best result from all layers. For more details on the linguistic and abstract concept information in different model layers, refer to Appendix E.2 and G.2.
Figure 1: Multilingual concept recognition accuracy (%) of LLaMA2-chat, Qwen-chat and BLOOMZ series, averaged across all value concepts. The performance of the three 7B-sized models are connected with dashed lines for performance comparison.
@@ 4.1 Experimental Setup
Our experiments involved three multilingual LLM families, including the LLaMA2-chat series (7B, 13B, 70B) (Touvron et al., 2023), Qwen-chat series (1B8, 7B, 14B) (Bai et al., 2023) and BLOOMZ series (560M, 1B7, 7B1) (Scao et al., 2022). Appendix D provides detailed language distributions of their pre-training data.
@@ 4.2 Q1: Do LLMs Encode Concepts Representing Human Values in Multiple Languages?
Figure 1 illustrates the multilingual concept recognition accuracy of the three LLM families, averaged across all value concepts. We observe that all three models achieve notable accuracy across all represented languages [^2] The model’s understanding in unrepresented languages may stem from cross-lingual transfer from other languages. Qwen’s technical report only mentions the inclusion of en and zh in its pre-training data. We conjecture the inclusion of 10 other languages (fr,es,pt,vi,ca,id,ja,ko,fi,hu) based on its significant performance in these languages. and even the smallest models surpass $\tau=65\%$ accuracy in them. These results confirm that LLMs effectively encode value concepts in a multilingual context. Moreover, Figure 1 shows a clear pattern that increasing model parameters and language resources notably improves concept recognition accuracy. This indicates that both expanding the model size and increasing the scale of language pre-training data contribute to a more precise encoding of value concepts.
Appendix E.1 compares the PCA-based method with the mean-based method outlined in §3.1. It reveals that both methods produce concept vectors of comparable precision, with the mean-based technique holding a slight edge. The consistent performance across various extraction techniques confirm the effectiveness of concept vectors in capturing conceptual information. Appendix E.2 indicates that middle layers are particularly adept at capturing abstract concept information. Appendix E.3 demonstrates that even a small number of training samples can effectively extract representations of value concepts in LLMs. For detailed results on each value concept and additional discussions, please refer to Appendix E.4 and E.5.
@@ 4.3.1 Trait 1: Inconsistency of Concept Representations between High- and Low-Resource Languages
The findings from Steck et al. (2024) suggest that a high average cosine similarity might raise concerns when dealing with unrelated representations. However, the results in Appendix G.1 indicate that, in our specific context, cosine similarity between concept vectors could reflect their genuine correlation. Additionally, Appendix G.2 presents cross-lingual consistency across different model layers, revealing that intermediate model layers exhibit greater consistency. For comprehensive results on each value concept and further discussions, please refer to Appendix G.3 and G.4.
Genetic Syntactic Geographic Phonological
D. C. D. C. D. C. D. C.
LLaMA2 -chat 7B -0.04 0.77 -0.12 0.63 -0.25 0.21 -0.03 -0.06
13B -0.17 0.53 -0.12 0.65 -0.17 0.35 0.09 0.24
70B -0.07 0.78 -0.12 0.66 -0.26 0.3 -0.0 0.01
Qwen -chat 1B8 0.06 0.42 0.07 0.32 -0.03 0.0 -0.02 0.05
7B 0.03 0.39 0.07 0.33 -0.04 0.04 -0.01 0.17
14B 0.01 0.42 0.01 0.5 -0.03 0.14 0.01 0.14
BLOOMZ 560M 0.20 0.43 0.13 0.55 -0.03 0.38 -0.12 -0.29
1B7 0.23 0.45 0.21 0.67 -0.01 0.43 -0.13 -0.28
7B1 0.16 0.36 0.09 0.52 -0.06 0.31 -0.11 -0.26
@@ 4.3.2 Trait 2: Linguistic Relationships Distortion due to the Imbalance of Language Data
To explore the correlation between cross-lingual concept consistency and linguistic similarity, following Qi et al. (2023), we used lang2vec [^4] https://github.com/antonisa/lang2vec to compute four types of linguistic similarity (genetic, syntactic, geographic, and phonological) between languages. We then calculated the Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs.
We employed two calculation methods to estimate the correlation. The first method directly computes the Pearson correlation on all language pairs (Direct), while the second starts by categorizing language pairs based on language resources. Subsequently, correlations are computed within different categories and averaged (Category). Please refer to Appendix F for details of the latter method.
@@ Appendix C Impact of Translation Quality
Figure 4: Multilingual concept recognition accuracy across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
@@ E.1 Extracting Concept Vectors based on PCA
E.2 Results across Model Layers
Figure 4 presents the concept recognition accuracy across different model layers. We observe that the most explicit concept vectors are encoded in the middle layers of the models. This suggests that the middle layers are more likely to encode explicit abstract value information, aligning with the findings of Li et al. (2023); Zou et al. (2023a).
E.3 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
@@ E.4 Complete Results
E.5 Multilingual Performance Reflects Multilinguality
@@ Appendix F Computing Pearson Correlation Coefficients Considering Differences in Language Resources
Figure 6: Cross-lingual similarity of concept vectors across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
@@ G.1 Cosine Similarity between Concept Vectors can Reflect Their Correlation
G.2 Results across Model Layers
Figure 6 illustrates the trends in cosine similarity across different model layers. We observe that the peak of cross-lingual consistency appears in the intermediate layers, with lower similarity near the input and output layers. This observation is consistent with previous research (Chi et al., 2021; Bhattacharya and Bojar, 2023), suggesting that middle layers of multilingual models encode a higher degree of language-independent information, while language-specific information is more prominent near the input and output layers.
G.3 Complete Results

### v3

@@ Abstract
Prior research has revealed that certain abstract concepts are linearly represented as directions in the representation space of LLMs, predominantly centered around English. In this paper, we extend this investigation to a multilingual context, with a specific focus on human values-related concepts (i.e., value concepts) due to their significance for AI safety. Through our comprehensive exploration covering 7 types of human values, 16 languages and 3 LLM series with distinct multilinguality (e.g., monolingual, bilingual and multilingual), we first empirically confirm the presence of value concepts within LLMs in a multilingual format. Further analysis on the cross-lingual characteristics of these concepts reveals 3 traits arising from language resource disparities: cross-lingual inconsistency, distorted linguistic relationships, and unidirectional cross-lingual transfer between high- and low-resource languages, all in terms of value concepts. Moreover, we validate the feasibility of cross-lingual control over value alignment capabilities of LLMs, leveraging the dominant language as a source language. Ultimately, recognizing the significant impact of LLMs’ multilinguality on our results, we consolidate our findings and provide prudent suggestions on the composition of multilingual data for LLMs pre-training.
@@
Shaoyang Xu1, Weilong Dong2, Zishan Guo2, Xinwei Wu2and Deyi Xiong2,1thanks:   Corresponding author 1School of New Media and Communication, Tianjin University, Tianjin, China 2College of Intelligence and Computing, Tianjin University, Tianjin, China {syxu, willowd, guozishan, wuxw2021, dyxiong}@tju.edu.cn
@@ 1 Introduction
Recent years have witnessed the emergence of large language models, such as ChatGPT (OpenAI, 2023a), GPT-4 (OpenAI, 2023b), and LLaMA2 (Touvron et al., 2023). These LLMs have shown powerful capabilities in natural language understanding and generation (Guo et al., 2023; Bang et al., 2023; Jiao et al., 2023; Liu et al., 2024). However, alongside with their prowess, LLMs present potential risks. Research has demonstrated that LLMs can generate responses containing toxic, untruthful, biased, and even illegal content (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023; Shen et al., 2023). Thus, aligning LLMs with human values (i.e., value alignment) is necessary for unleashing their potential safely.
@@
However, existing studies on concept representations in LLMs have primarily focused on English (Zou et al., 2023a), leaving multilingual concepts unexplored. Our work is the first to explore multilingual concepts in LLMs, emphasizing human values-related concepts to advance multilingual AI safety and utility. The primary research questions we aim to answer are as follows: (Q1) Do LLMs encode concepts representing human values in multiple languages? (Q2) To what extent are these concepts consistent and transferable across different languages? (Q3) Whether LLMs trained with different distributions of multilingual data exhibit distinct multilinguality in these concepts? (Q4) Is Value Alignment of LLMs Controllable across Languages? To address these questions, we propose a framework consisting of 5 components: extracting multilingual concept vectors from LLMs (§3.1) and evaluating their correlation with the corresponding concepts (concept recognition task in §3.2) to answer Q1; computing cross-lingual similarity of concept vectors (§3.3) and performing cross-lingual concept recognition (§3.4) to answer Q2 and Q3; and manipulating model behavior cross-lingually via concept vectors (§5) to answer Q4.
@@
Our analysis covers 7 concepts related to human values: commonsense morality, deontology, utilitarianism, fairness, truthfulness, toxicity and harmfulness, given their significance for AI safety (Hendrycks et al., 2021; Bai et al., 2022; Askell et al., 2021; Touvron et al., 2023; Yu et al., 2024; Shen et al., 2023; Guo et al., 2023). To ensure the breadth and reliability of our findings, we have selected these 7 concepts for their diverse definitions and ethical attributes (Vida et al., 2023). Throughout this paper, we collectively refer to them as “value concepts” to reflect their diversity and keep consistent with existing AI alignment research (Bai et al., 2022; Askell et al., 2021; Hendrycks et al., 2021). For comprehensive definitions, ethical backgrounds and examples of these value concepts, please refer to Appendix A.
@@
In addition to diverse human values, our experiments involve 16 languages [^1] We recognize that linguistic diversity can foster cultural variations, potentially resulting in diverse interpretations of the same value from different cultural backgrounds (Hershcovich et al., 2022; Hämmerl et al., 2023). For example, regarding deontology, some cultures prioritize individual responsibility while others emphasize social obligations (Cao et al., 2023; Hofstede, 1984). However, our work focuses on the multilingual representations of value concepts within LLMs and their universal cross-lingual patterns, leaving the exploration on cultural divergences in human values for our future research. and 3 LLM families with different multilinguality. Specifically, we categorize the multilinguality of these 3 LLM families based on language distributions in their pre-training data into 3 groups: English-dominated LLMs (LLaMA2-chat series in our experiments), Chinese & English-dominated LLMs (i.e., Qwen-chat series), and LLMs with more balanced multilinguality (i.e., BLOOMZ series). Appendix D provides detailed language distributions of their pre-training data.
@@ 2 Related Work
Representation Engineering (RepE) introduced by Zou et al. (2023a) extracts abstract concepts as vectors from LLMs using positive and negative samples that describe specific concepts. The effectiveness of these vectors has been validated across dimensions such as correlation and manipulation. Specifically, correlation experiments have assessed the predictive power of the extracted vectors to classify out-of-distribution data as positive or negative, while manipulation experiments have evaluated the vectors’ ability to control LLMs’ behavior by adding or subtracting them from the hidden states (Liu et al., 2023; Leong et al., 2023; Wang and Shu, 2023; Wu et al., 2024; Dong et al., 2024). While previous research has primarily focused on English, we pioneer the extension of RepE into a multilingual context, exploring multilingual concepts within LLMs through concept extraction, correlation, and manipulation experiments, all conducted in a multilingual or cross-lingual manner.
@@
Despite their remarkable capabilities, LLMs present potential risks (Cui et al., 2024; Wang et al., 2023; Huang et al., 2023; Shi and Xiong, 2024; Huang and Xiong, 2024). Research has indicated that LLMs exhibit even greater vulnerability in multilingual scenarios, as evidenced by multilingual jailbreak challenges (Deng et al., 2023; Shen et al., 2024; Yong et al., 2023). In this work, we explore the cross-lingual characteristics of safety-related concepts in LLMs’ representation spaces, shedding light on the mechanisms underlying multilingual vulnerability.
@@ 3.2 Recognizing Multilingual Concepts
Note that each layer has a recognition accuracy, using the concept vector of that layer. Unless specified otherwise, we report the best accuracy.
Figure 1: Multilingual concept recognition accuracy (%) of LLaMA2-chat, Qwen-chat and BLOOMZ series, averaged across all value concepts. The performance of the three 7B-sized models are connected with dashed lines for performance comparison. “Represented languages” refer to the languages present in the pre-training corpus.
Figure 2: (a) Multilingual concept recognition accuracy across different model layers. (b) Cross-lingual similarity of concept vectors across different model layers. Results are averaged across languages included both in LLaMA2-chat and BLOOMZ series’ pre-training data, as well as across all human values.
@@ 4.1 Experimental Setup
Our experiments involved three multilingual LLM families, including the LLaMA2-chat series (7B, 13B, 70B) (Touvron et al., 2023), Qwen-chat series (1B8, 7B, 14B) (Bai et al., 2023) and BLOOMZ series (560M, 1B7, 7B1) (Scao et al., 2022). Appendix D provides detailed language distributions of their pre-training data. Notably, not all selected languages are included in the pre-training data of these model families. Specifically, both LLaMA2 and BLOOMZ cover 12 of these languages, though their selections do not fully overlap. In contrast, Qwen’s technical report only mentions the inclusion of English and Chinese. For the multilingual concept recognition task, we consider all 16 languages, regardless of the model series, while other tasks explore only the languages covered in the pre-training data.
@@ 4.2 Q1: Do LLMs Encode Concepts Representing Human Values in Multiple Languages?
Figure 1 illustrates the multilingual concept recognition accuracy of the three LLM families, averaged across all value concepts. We first observe that all three models achieve notable accuracy across all represented languages and even the smallest models surpass $\tau=65\%$ accuracy in them. It’s important to note that the accuracy of $65\%$ is a conservative statistic and represents a lower bound, derived from the smallest model (BLOOMZ-560M) on the poorest-performing language (ny, accounting for only 0.00007% in pre-training data). However, results from larger models are significantly higher. For example, BLOOMZ-7B1 achieves accuracy exceeding 81% on the majority of seen languages (10 out of 12). In addition to BLOOMZ-7B1, other model families with equivalent model sizes also demonstrate similarly high performance. Overall, these results confirm that LLMs effectively encode value concepts in a multilingual context.
We also observe a certain level of recognition accuracy in some unrepresented languages. We conjecture that the ability of models in capturing these languages may stem from cross-lingual transfer from other languages. Additionally, as mentioned in Section 4.1, Qwen’s technical report only mentions the inclusion of en and zh in its pre-training data. We conjecture the inclusion of 10 other languages (fr,es,pt,vi,ca,id,ja,ko,fi,hu) based on its significant performance in these languages.
Although previous results represent the best performance across all layers, Figure 2(a) presents the concept recognition accuracy across different model layers. We observe that middle layers encode more abstract information related to human values, aligning with the findings of Li et al. (2023).
Appendix E.1 compares the PCA-based method with the mean-based method outlined in §3.1. It reveals that both methods produce concept vectors of comparable precision, with the mean-based technique holding a slight edge. The consistent performance across various extraction techniques confirm the effectiveness of concept vectors in capturing conceptual information. Appendix E.2 demonstrates that even a small number of training samples can effectively extract representations of value concepts in LLMs. For detailed results on each value concept and additional discussions, please refer to Appendix E.3 and E.4.
Genetic Syntactic Geographic Phonological
D. C. D. C. D. C. D. C.
LLaMA2 -chat 7B -0.04 0.77 -0.12 0.63 -0.25 0.21 -0.03 -0.06
13B -0.17 0.53 -0.12 0.65 -0.17 0.35 0.09 0.24
70B -0.07 0.78 -0.12 0.66 -0.26 0.30 0.00 0.01
Qwen -chat 1B8 0.06 0.42 0.07 0.32 -0.03 0.00 -0.02 0.05
7B 0.03 0.39 0.07 0.33 -0.04 0.04 -0.01 0.17
14B 0.01 0.42 0.01 0.50 -0.03 0.14 0.01 0.14
BLOOMZ 560M 0.20 0.43 0.13 0.55 -0.03 0.38 -0.12 -0.29
1B7 0.23 0.45 0.21 0.67 -0.01 0.43 -0.13 -0.28
7B1 0.16 0.36 0.09 0.52 -0.06 0.31 -0.11 -0.26
@@ 4.3.1 Trait 1: Inconsistency of Concept Representations between High- and Low-Resource Languages
Additionally, Figure 2(b) illustrates the trends in cosine similarity across different model layers. We observe that the peak of cross-lingual consistency appears in the intermediate layers, with lower similarity near the input and output layers. This observation is consistent with previous research (Chi et al., 2021; Bhattacharya and Bojar, 2023), suggesting that middle layers of multilingual models encode a higher degree of language-independent information, while language-specific information is more prominent near the input and output layers.
The findings from Steck et al. (2024) suggest that a high average cosine similarity might raise concerns when dealing with unrelated representations. However, the results in Appendix G.1 indicate that, in our specific context, cosine similarity between concept vectors could reflect their genuine correlation. For comprehensive results on each value concept and further discussions, please refer to Appendix G.2 and G.3.
@@ 4.3.2 Trait 2: Linguistic Relationships Distortion due to the Imbalance of Language Data
Figure 3 also suggests that LLMs may learn linguistic correlations between languages and reflect them in cross-lingual concept consistency. Regarding BLOOMZ-7B1, although the cross-lingual consistency between the low-resource languages ta, te and other languages is low, the consistency between these two languages is very high because they both belong to the Dravidian language family. A similar pattern is observed for sw and ny, both of which are from the Niger-Congo family. [^3] This trend also applies to LLaMA2-chat-7B, where the cross-lingual consistency between en and fr, es, pt, ca is higher because they all belong to the Indo-European language family. From this observation, we hypothesize that cross-lingual concept consistency may be influenced by both the amount of language resources and linguistic relationships between languages. In this section, we further explore this phenomenon, specifically investigating to what extent cross-lingual concept consistency reflects natural linguistic relationships between languages and how language resources affect their correlation.
To explore the correlation between cross-lingual concept consistency and linguistic similarity, following Qi et al. (2023), we used lang2vec [^4] https://github.com/antonisa/lang2vec to compute four types of linguistic similarity (genetic, syntactic, geographic, and phonological) between languages. We then calculated the Pearson correlation between cross-lingual concept consistency and linguistic similarity for all language pairs. We employed two calculation methods to estimate the correlation. The first method directly computes the Pearson correlation on all language pairs (Direct), while the second starts by categorizing language pairs based on language resources. Subsequently, correlations are computed within different categories and averaged (Category). Such categorization aims to mitigate the influence of language resources. Please refer to Appendix F for details of the latter method.
@@ 7 Conclusion
Acknowledgements
The present research was supported by the National Key Research and Development Program of China (Grant No. 2023YFE0116400). We would like to thank the anonymous reviewers for their insightful comments.
@@ E.1 Extracting Concept Vectors based on PCA
E.2 Varying the Size of $\mathcal{T}_{c}^{\textnormal{train}}$
@@ E.3 Complete Results
E.4 Multilingual Performance Reflects Multilinguality
@@ G.1 Cosine Similarity between Concept Vectors can Reflect Their Correlation
G.2 Complete Results